solve by the substitution method
9x+4y= -24
x=20-8y

Answer by CharlesG2(834)

solve by the substitution method
9x+4y= -24
x=20-8y
9 * (20 - 8y) + 4y = -24 (plugged 2nd into first)
180 - 72y + 4y = -24
180 - 68y = -24
-68y = -204
68y = 204
y = 3
x = 20 - 8(3)
x = 20 - 24 = -4